# Representation Shattering in Transformers: A Synthetic Study with Knowledge Editing

### v3

@@ 3.4 Representation shattering
where $||D||_{F}$ is the Frobenius norm of $D$, $D_{\varnothing}$ the pairwise distance matrix of the entities computed using the unedited model, and $D_{*}$ is the pairwise distance matrix computed using the edited model. The distance between entities is computed by measuring the euclidean distance between the representation vector of each entity.
@@ 4 Uncovering Representation Shattering
Test type (a) Unedited model (b) Corrective edits (c) $\langle\Delta\text{Acc.}\rangle$. ⟩ for Counterfactual edits Cyclic Order Acc. Sub-Graph $\langle\Delta\text{Acc.}\rangle$. ⟩ $d=1$ $d=2$ $d=3$ $d=4$ Direct recall I 98.34 Edit -21.95 -01.49 -67.01 -77.07 -77.94 II 93.71 Retain -22.64 -01.91 -66.70 -75.49 -75.42 III 99.37 Test -21.83 -01.75 -67.00 -76.12 -77.90 Logical inference I 98.16 Edit -22.24 -01.44 -67.22 -77.14 -78.02 II 93.95 Retain -22.50 -01.83 -66.88 -75.67 -75.67 III 99.40 Test -22.03 -01.80 -67.31 -76.27 -78.23 Compositional inference I 88.15 Edit -29.60 -05.32 -73.15 -80.35 -80.63 II 79.31 Retain -31.92 -05.32 -71.21 -78.70 -78.87 III 93.50 Test -31.70 -06.69 -74.88 -81.38 -80.62
Table 1: The direct recall, logical inference, and compositional inference accuracies before and after KE. Results are for ROME; see Appx. F.1.1 for other methods. (a) The performance of our model (before editing) across the three cyclic orders (I, II, and III). Not only does our model perform well on direct recall, but it also generalizes to both logical and compositional inference tasks. This suggests that the model’s internal representations extend beyond simple memorization and capture the underlying global structure that relates entities. (b) Changes in model accuracy after applying corrective knowledge edits. Each $\langle\Delta\text{Acc.}\rangle$. ⟩ result is averaged across multiple edits, and each row labeled edit/retain/test is averaged across each of the cyclic orders taking turns, i.e., playing the roles of the edit, retain, and test sub-graphs. We find that corrective knowledge edits negatively affect the model’s accuracy both on related and unrelated facts. These results align with the findings on LLMs (Gu et al., 2024; Gupta et al., 2024a). (c) $\langle\Delta\text{Acc.}\rangle$. ⟩ for edit, retain, and test sub-graphs after applying counterfactual edits. Intentionally introducing inconsistencies into the model’s knowledge via counterfactual KE can significantly degrade model capabilities. Furthermore, the greater the induced inconsistency (scaling the counterfactual edit distance $d$ from 1-4), the more severe the resulting performance degradation.
We study knowledge editing methods like ROME (Meng et al., 2022a), MEMIT (Meng et al., 2022b), PMET (Li et al., 2024), and AlphaEdit (Fang et al., 2024) in this work. While in the main paper we primarily present results with ROME (see Appx. D.1 for a short primer), we provide results with other methods in Appx. F.1.1 and Appx. F.1.2. We perform two different types of edits: corrective edits and counterfactual edits. Corrective edits are applied to facts which the model recalls incorrectly after training. A counterfactual edit introduces a new fact, i.e., it changes fact $(x_{i},r,x_{j})$ to fact $(x_{i},r,x_{k})$ where $x_{j}\neq x_{k}$. Such an edit introduces inconsistencies in the knowledge graph.
@@
4.1 Evaluating the effects of knowledge editing
@@
Figure 4: Transformers learn representations that mirror the geometry of the underlying data. (a) The representations—or output of the second attention layer—for the input $x\vec{r}$ for different entities $x$ and fixed relation $\vec{r}$ are visualized using Isomap. The model learns the cyclic ordering to represent all the facts. For visualizations of representations at various other model components and relation prompts, please see Appx. F.3 and Appx. F.4. (b) To improve the visual fidelity of the projected representations when comparing representations of post-edit models to the unedited model, we construct Isomap neighborhood graphs using the outputs of the Transformer. For more details, please see Appx. E.
4.2 Transformer representations capture the geometry of the data
The model achieves high compositional and logical inference accuracies before knowledge editing, indicating that it captures the global structure of the data and does not merely memorize all the facts seen during training. We see this reflected in the internal representation of the model (output of the second attention layer), which we visualize using Isomap (Tenenbaum et al., 2000)—a non-linear dimensionality reduction method that uses multi-dimensional scaling with distances computed using a local neighborhood graph. In Fig. 4a, we plot the evolution of the Isomap embedding—of the internal representation for the input with one entity and relation ($x\vec{r}$)—over the course of training. The different data points correspond to different values of the entity $x$, for a fixed relation $\vec{r}$ and the points in the plot are colored by the cyclic ordering. We see that the representation manifold resembles the cyclic ordering of the entities, particularly towards the end of training. In addition, we confirm that the model correctly internalizes the independence of the cyclic orders I, II, and III in its representations (see Appx. F.2 for a visualization).
4.3 Corrective knowledge edits shatter the representation geometry
@@
Figure 5: Representation shattering strongly correlates with a degradation in accuracy. (a) We plot the change in direct recall, logical inference, and compositional inference accuracies for different edits models, edited on different facts. We find all 3 accuracies to be strongly correlated. We select 3 edited models that span the range of accuracies, which is denoted by and  in the plot. (b) We plot the representations using a variant of Isomap (see Appx. E) with the entities colored by the cyclic order. We observe a clear trend where larger drop in accuracy directly correlates with a greater degree of representation shattering, i.e., the geometric structure of the data is destroyed after the edit. (c) We plot the mean drop in accuracy against the representation shattering metric $R(D_{*})$ as defined in Eq. 1. Greater representation shattering is strongly correlated with more severe accuracy degradation ($r^{2}=0.905$).
@@
We visualize the representations of 3 different models using the techniques described in Sec. 4.2. The 3 models are obtained after applying 3 different edits and are selected to have high (), intermediate (), and low () direct recall accuracies. In Fig. 5, we observe that the model with the highest accuracy () has a representation that preserves the geometry of the data after the edit. However, as the model accuracy decreases, the representations also display a greater degree of distortion, no longer capturing the geometry of the data; in other words, the model is affected by representation shattering. Beyond visual inspection, this trend is also quantified in Fig. 5c, which shows a strong negative relationship between the distortion metric $R(D_{*})$ (Eq. 1) and model accuracy ($r^{2}=0.905$).
4.4 How do different counterfactual edits change the extent of shattering?
@@
Table 2: Mean $R(D_{*})$ for counterfactual edits, averaged across each sub-graph type. We observe higher degrees of representation shattering for greater counterfactual edit distances ($d$). Results are for ROME; other methods also reproduce this relationship (Appx. F.1.2).
@@
Figure 6: Counterfactual edits with larger edit distance result in larger drop in accuracy and greater degree of representation shattering. We apply counterfactual knowledge edits to overwrite a correctly learned fact (1154.I_C1=567) with inconsistent counterfactual associations. We then plot the accuracy after the counterfactual edit for different edit distances and the corresponding low-dimensional embedding of the representation obtained using a modified version of Isomap (see Appx. E). The numerical quantity in the upper right of each manifold visualization is the $R(D_{*})$ value measuring the degree of representation shattering with respect to the manifold of the unedited model. Both visually and numerically, we find that a counterfactual edit with larger edit distance requires a significant distortion to the representation geometry to learn the new fact.
@@
Counterfactual editing, wherein ones adds new facts that were unseen during training, is commonly used for evaluating KE protocols (Meng et al., 2022a; 2023; Gupta et al., 2024a; Hoelscher-Obermaier et al., 2023). We consider 25 different counterfactual edits corresponding to every single counterfactual edit distance, where the counterfactual edit distance (or CE distance) is the distance between the entity in the old fact and new fact as measured in the cyclic order. Fig. 3 illustrates an example where the counterfactual edit has an edit distance of 1. In Fig. 6, we see that increasing the distance of the counterfactual edit results in a drop in accuracy and an increasing in the extent of shattering. This relationship is numerically supported by $R(D_{*})$, as shown in Tab. 2: shattering increases as counterfactual edit distance increases. In other words, when a new fact changes one entity to another, the extent of shattering increases as the distance between the old and new entity increases. As a naturalistic parallel, if the entities are different months, accuracy is higher when we edit “December” to “November” as opposed to “July”.
4.5 Representation shattering in LLMs
@@
Finally, we investigate whether our findings generalize to large Transformers trained on naturalistic data. We consider concepts with a cyclic order, in particular months of the year, and apply counterfactual edits to Llama 3 (Grattafiori et al., 2024) and Mamba S4 (Gu & Dao, 2023) (see Appx. G.1). We additionally explore non-cyclic geometries, specifically tree-structured concepts, and their representation shattering in Appx. G.3.
@@
In Fig. 7a, we examine how the reasoning performance of Llama 3 8B Instruct is affected by KE in relation to the counterfactual edit distance. We find that as the edit distance increases, there is a gradual decline in the model’s reasoning accuracy. Furthermore, Fig. 7b shows the latent representations for the 12 months, extracted from an intermediate layer post-editing. As the edit distance is varied from 1 to 3, the observed degree of representation shattering also increases. This result demonstrates that our insights from synthetic data, i.e. the representation shattering hypothesis, can generalize to larger models trained on naturalistic data.
Figure 7: Representation shattering also occurs in a real LLM (Llama 3 8B Instruct). We apply KE (MEMIT) with counterfactual prompts to modify the cyclic order of calendar months in the Llama 3 8B Instruct model. After editing, we evaluate the model’s ability to perform reasoning on the MMLU-Redux benchmark and visualize the model’s representations for each month. (a) As the counterfactual edit distance grows, the model’s reasoning abilities gradually decrease. (b) Coinciding with the degredation in reasoning performance, the gradual shattering of the ring structure in the model’s intermediate representations occurs as the counterfactual edit distance is increased (activations were collected at the layer 19 output; see Appx. G.2 for other extraction points). For similar results with Mamba S4 (Gu & Dao, 2023), refer to Appx. G.1. For experiments with concepts organized in a non-circular geometry, refer to Appx. G.3.
@@ 5 Conclusion
In this work, we introduced a synthetic framework to analyze the side effects of knowledge editing in transformers, identifying “representation shattering” as a key factor behind performance degradation. Specifically, we show preserving representational structures underlying a model’s knowledge is crucial to avoiding negative consequences of knowledge editing: distortion of such structures impacts a model’s broader capabilities. To arrive at this hypothesis, we design a controlled framework that allows investigations into models modified by knowledge editing protocols, offering clear representation-level explanations for why knowledge editing can harms models’ broader capabilities that generalize to real-world models like Llama 3. While the use of simplified tasks and models can limit the scope of our conclusions, since larger, more complex real-world models may exhibit additional dynamics that our framework does not capture, we do believe that testing knowledge editing protocols on setups similar to our synthetic, knowledge graph one will significantly aid design of better editing protocols. We claim failing even such simple, albeit systematically defined settings, likely implies the editing protocol should not be readily trusted or applied at scale.

### v5

@@ 3.4 Representation Shattering
where $||D||_{F}$ is the Frobenius norm of $D$, $D_{\varnothing}$ is the pairwise distance matrix of the entities computed using the unedited model, and $D_{*}$ is the pairwise distance matrix computed using the edited model. $D_{*}$ and $D_{\varnothing}$ are both $n\times n$ matrices over $n$ entities, and entry $D_{ij}$ is the Euclidean distance between the representation vectors of $x_{i}$ and $x_{j}$. The matrices are computed over only entity tokens (relation tokens are excluded).
Note that $R(D_{*})$ is permutation-sensitive by design: a zero value indicates that every entity token preserves its location in representation space. Isomorphic geometries (e.g., wherein representations of two entity tokens swap positions in the manifold) still yield nonzero $R(D_{*})$ values, because the relative positions of entities (as identified by their respective pre-assigned tokens) must have been distorted to achieve the edit(s).
@@ 4 Uncovering Representation Shattering
(a) Unedited (b) Corrective (c) Counterfactual Cyclic Order Acc. Sub- Graph $\langle\Delta\text{Acc.}\rangle$. ⟩ Distance-$d$ Edit $\langle\Delta\text{Acc.}\rangle$. ⟩ $d$=1 $d$=2 $d$=3 $d$=4 Direct Recall I 98.34 Edit -21.95 -01.49 -67.01 -77.07 -77.94 II 93.71 Retain -22.64 -01.91 -66.70 -75.49 -75.42 III 99.37 Test -21.83 -01.75 -67.00 -76.12 -77.90 Logi. Infer. I 98.16 Edit -22.24 -01.44 -67.22 -77.14 -78.02 II 93.95 Retain -22.50 -01.83 -66.88 -75.67 -75.67 III 99.40 Test -22.03 -01.80 -67.31 -76.27 -78.23 Comp. Infer. I 88.15 Edit -29.60 -05.32 -73.15 -80.35 -80.63 II 79.31 Retain -31.92 -05.32 -71.21 -78.70 -78.87 III 93.50 Test -31.70 -06.69 -74.88 -81.38 -80.62
Table 1: The direct recall, logical inference, and compositional inference accuracies before and after KE. Results are for ROME; see Appx. G.1.1 for other methods. (a) The performance of our model (before editing) across the three cyclic orders (I, II, and III). Not only does our model perform well on direct recall, but it also generalizes to both logical and compositional inference tasks. This suggests that the model’s internal representations extend beyond simple memorization and capture the underlying global structure that relates entities. (b) Changes in model accuracy after applying corrective knowledge edits. Each $\langle\Delta\text{Acc.}\rangle$. ⟩ result is averaged across multiple edits, and each row labeled edit/retain/test is averaged across each of the cyclic orders taking turns, i.e., playing the roles of the edit, retain, and test sub-graphs. We find that corrective knowledge edits negatively affect the model’s accuracy both on related and unrelated facts. These results align with the findings on LLMs (Gu et al., 2024; Gupta et al., 2024a). (c) $\langle\Delta\text{Acc.}\rangle$. ⟩ for edit, retain, and test sub-graphs after applying counterfactual edits. Intentionally introducing inconsistencies into the model’s knowledge via counterfactual KE can significantly degrade model capabilities. Furthermore, the greater the induced inconsistency (scaling the counterfactual edit distance $d$ from 1-4), the more severe the resulting performance degradation.
We study knowledge editing methods like ROME (Meng et al., 2022a), MEMIT (Meng et al., 2022b), PMET (Li et al., 2024), and AlphaEdit (Fang et al., 2024) in this work. While in the main paper we primarily present results with ROME (see Appx. E.1 for a short primer), we provide results with other methods in Appx. G.1.1 and Appx. G.1.2. We also verify the assumptions made by these methods in Appx. G.2.
We perform two different types of edits: corrective edits and counterfactual edits. Corrective edits are applied to facts which the model recalls incorrectly after training. A counterfactual edit introduces a new fact, i.e., it changes fact $(x_{i},r,x_{j})$ to fact $(x_{i},r,x_{k})$ where $x_{j}\neq x_{k}$. Such an edit introduces inconsistencies in the knowledge graph.
@@
4.1 Knowledge Editing can Degrade Model Accuracy
@@
Figure 4: Transformers learn representations that mirror the geometry of the underlying data. (a) The representations—or output of the second attention layer—for the input $x\vec{r}$ for different entities $x$ and fixed relation $\vec{r}$ are visualized using Isomap. The model learns the cyclic ordering to represent all the facts. For visualizations of representations at various other model components and relation prompts, please see Appx. G.4 and Appx. G.5. (b) To improve the visual fidelity of the projected representations when comparing post-edit models to the unedited model, we pre-process the Isomap neighborhood graphs using the outputs of the Transformer model. For more details, see Appx. F.
4.2 Model Representations Capture Data Geometry
The model achieves high compositional and logical inference accuracies before knowledge editing, indicating that it captures the global structure of the data and does not merely memorize all the facts seen during training. We see this reflected in the internal representation of the model (output of the second attention layer), which we visualize using Isomap (Tenenbaum et al., 2000). Isomap preserves geodesic distances along the manifold, which is essential for visualizing non-linear geometries—e.g., rings and torii (Chaudhuri et al., 2019; Khona & Fiete, 2022). Meanwhile, PCA, being linear, can collapse or stretch such structures. This makes us prefer Isomap for our analysis, since faithful projections of the original topology turn out to be critical for interpreting how KE affects global geometry. We use PCA as well when linearity suffices (see Appx. G.2.2 and Appx. G.3).
In Fig. 4a, we plot how the Isomap embeddings of the internal representations for the input with one entity and relation ($x\vec{r}$) evolves over the course of training. The different data points correspond to different values of the entity $x$, for a fixed relation $\vec{r}$ and the points in the plot are colored by the cyclic ordering. We see that the representation manifold resembles the cyclic ordering of the entities, particularly towards the end of training. In addition, we confirm that the model correctly internalizes the independence of the cyclic orders I, II, and III in its representations (see Appx. G.3 for a visualization).
4.3 Corrective KE Shatters Representation Geometry
@@
Figure 5: Representation shattering strongly correlates with a degradation in accuracy. (a) We plot the change in direct recall, logical inference, and compositional inference accuracies for different edits models, edited on different facts. We find all 3 accuracies to be strongly correlated. We select 3 edited models that span the range of accuracies, which is denoted by and  in the plot. (b) We plot the representations using a variant of Isomap (see Appx. F) with the entities colored by the cyclic order. We observe a clear trend where larger drop in accuracy directly correlates with a greater degree of representation shattering, i.e., the geometric structure of the data is destroyed after the edit. (c) We plot the mean drop in accuracy against the representation shattering metric $R(D_{*})$ as defined in Eq. 1. Greater representation shattering is strongly correlated with more severe accuracy degradation ($r^{2}=0.905$).
@@
We visualize the representations of 3 different models using the techniques described in Sec. 4.2 and Fig. 4b. The 3 models are obtained after applying 3 different edits and are selected to have high (), intermediate (), and low () direct recall accuracies. In Fig. 5, we observe that the model with the highest accuracy () has a representation that preserves the geometry of the data after the edit. However, as the model accuracy decreases, the representations also display a greater degree of distortion, no longer capturing the geometry of the data; in other words, the model is affected by representation shattering. Beyond visual inspection, this trend is also quantified in Fig. 5c, which shows a strong negative relationship between the distortion metric $R(D_{*})$ (Eq. 1) and model accuracy ($r^{2}=0.905$).
4.4 Shattering Scales with Counterfactual Edit Distance
@@
Table 2: Mean $R(D_{*})$ for counterfactual edits, averaged across each sub-graph type. We observe higher degrees of representation shattering for greater counterfactual edit distances ($d$). Results are for ROME; other methods also reproduce this relationship (Appx. G.1.2).
@@
Figure 6: Counterfactual edits with larger edit distance result in larger drop in accuracy and greater degree of representation shattering. We apply counterfactual knowledge edits to overwrite a correctly learned fact (1154.I_C1=567) with inconsistent counterfactual associations. We then plot the accuracy after the counterfactual edit for different edit distances and the corresponding low-dimensional embedding of the representation obtained using a modified version of Isomap (see Appx. F). The numerical quantity in the upper right of each manifold visualization is the $R(D_{*})$ value measuring the degree of representation shattering with respect to the manifold of the unedited model. Both visually and numerically, we find that a counterfactual edit with larger edit distance requires a significant distortion to the representation geometry to learn the new fact.
@@
Counterfactual editing, wherein ones adds new facts that were unseen during training, is commonly used for evaluating KE protocols (Meng et al., 2022a, 2023; Gupta et al., 2024a; Hoelscher-Obermaier et al., 2023). We consider 25 different counterfactual edits corresponding to every single counterfactual edit distance, where the counterfactual edit distance (or CE distance) is the distance between the entity in the old fact and new fact as measured in the cyclic order. Fig. 3 illustrates an example where the counterfactual edit has an edit distance of 1. In Fig. 6, we see that increasing the distance of the counterfactual edit results in a drop in accuracy and an increasing in the extent of shattering. This relationship is numerically supported by $R(D_{*})$, as shown in Tab. 2: shattering increases as counterfactual edit distance increases. In other words, when a new fact changes one entity to another, the extent of shattering increases as the distance between the old and new entity increases. As a naturalistic parallel, if the entities are different months, accuracy is higher when we edit “December” to “November” as opposed to “July”.
Summarizing these experiments, we find that for a given subject entity, edits with larger CE distances imply greater displacement in the representation manifold and higher $R(D_{*})$—in reverse, higher $R(D_{*})$ implies a larger CE distance. This manipulation approximates an intervention on representation geometry itself, and this offers compelling evidence for a causal link between shattering and performance loss. This is consistent with our mechanistic explanation, though we do not claim formal proof.
4.5 Representation Shattering Generalizes to LLMs
@@
Finally, we investigate whether our findings generalize to large Transformers trained on naturalistic data. We consider concepts with a cyclic order, in particular months of the year, and apply counterfactual edits to Llama 3 (Grattafiori et al., 2024) and Mamba S4 (Gu & Dao, 2023) (see Appx. H.1). We additionally explore non-cyclic geometries, specifically tree-structured concepts, in Appx. H.3.
@@
Figure 7: Representation shattering also occurs in a real LLM (Llama 3 8B Instruct). We apply KE (MEMIT) with counterfactual prompts to modify the cyclic order of calendar months in the Llama 3 8B Instruct model. After editing, we evaluate the model’s ability to perform reasoning on the MMLU-Redux benchmark and visualize the model’s representations for each month. (a) As the counterfactual edit distance grows, the model’s reasoning abilities gradually decrease. (b) Coinciding with the degredation in reasoning performance, the gradual shattering of the ring structure in the model’s intermediate representations occurs as the counterfactual edit distance is increased (activations were collected at the layer 19 output; see Appx. H.2 for other extraction points). For similar results with Mamba S4 (Gu & Dao, 2023), refer to Appx. H.1. For experiments with concepts organized in a non-circular geometry, refer to Appx. H.3.
In Fig. 7a, we examine how the reasoning performance of Llama 3 8B Instruct is affected by KE in relation to the counterfactual edit distance. We find that as the edit distance increases, there is a gradual decline in the model’s reasoning accuracy. Furthermore, Fig. 7b shows the latent representations for the 12 months, extracted from an intermediate layer post-editing. As the edit distance is varied from 1 to 3, the degree of representation shattering also increases. This result demonstrates that our insights from synthetic data, i.e. the representation shattering hypothesis, can generalize to larger models trained on naturalistic data.
@@ 5 Conclusion
In this work, we introduced a synthetic framework to analyze the side effects of knowledge editing in transformers, identifying “representation shattering” as a key factor behind performance degradation. Specifically, we show preserving representational structures underlying a model’s knowledge is crucial to avoiding negative consequences of knowledge editing: distortion of such structures impacts a model’s broader capabilities. To arrive at this hypothesis, we design a controlled framework that allows investigations into models modified by knowledge editing protocols, offering clear representation-level explanations for why knowledge editing can harms models’ broader capabilities that generalize to real-world models like Llama 3. While the use of simplified tasks and models can limit the scope of our conclusions, since larger, more complex real-world models may exhibit additional dynamics that our framework does not capture, we believe that testing knowledge editing protocols on setups similar to our synthetic knowledge graph will significantly aid the process of designing better editing protocols. Failing even such simple, albeit systematically defined settings, likely implies the editing protocol should not be readily trusted or applied at scale.
Additionally, our study aligns with a broader theoretical framework: (1) Transformers store factual associations in key-value pairs within MLP layers (Geva et al., 2020; Meng et al., 2022a), corroborated by our synthetic model analysis. (2) Parameter sharing and superposition cluster unrelated facts in overlapping subspaces (Geva et al., 2020; Henighan et al., 2023), making them vulnerable to unintentional interference. (3) Entities and relations often form structured manifolds (e.g., cycles, hierarchies), which aid compositional inference (Engels et al., 2024). (4) KE methods (ROME, MEMIT, etc.) enact local weight updates that deform these manifolds, causing representation shattering for unedited facts residing in shared sub-regions. Points (1)-(4) together lead us to believe that the fragility of KE arises from the entangled, compressed nature of factual storage, rather than just from the task of knowledge retention in and of itself.
More broadly, our study suggests that mitigating the pitfalls of KE may require fundamentally rethinking the popular “localize-then-edit” paradigm. Indeed, the field has already begun moving in this direction: promising methods include Retrieval Augmentation (Lewis et al., 2020; Borgeaud et al., 2022), Lifelong Model Editing (Wang et al., 2024), and Synthetic Document Finetuning (Wang et al., 2025). We hope that our work can help guide the design of further improvements to KE approaches.
Impact Statement
This paper introduces a scientific framework to analyze the phenomenon of representation shattering in Transformer models during Knowledge Editing. We believe the insights uncovered in this work will form a foundational step toward developing more robust and interpretable knowledge editing techniques, ultimately advancing the reliability of AI systems for societal benefit.